A 4.0 μF capacitor is put across a 12 V battery. After a long time (~20 time constants), the capacitor is disconnected and placed in series through an open switch, with a 250 Ω resistor.

(a) Determine the charge on the capacitor before it is discharged. C

(b) What is the initial current through the resistor when the switch is closed? A

(c) At what time will the current reach 37% of its initial value? s
Answer #1

(a) Ne = 12 Vedt, Q = Cue Q = (yef) (12) = 4846 Or 4881086 tot 1242 = 0.0484 120 - 0.0484 250st T= RC = (250.2) (45) = 18153